# Zihao Zhang, Guodong Li, and Sihuang Hu

### v1

@@ 3.1 The lower bound on avg-min repair bandwidth
Recall the definitions of set $\mathcal{R}(\mathbf{M})$ in (5) and set $\mathcal{M}_{i}$ in (6). According to $d(\mathbf{M})=k+1$, we have
$\operatorname{BW}(\mathbf{M})=\sum_{j\in[n]}\textnormal{rank}\left(\mathbf{M}\mathbf{H}_{j}\right)-2=k+|\mathcal{R}(\mathbf{M})|,$
which means that for a node $C_{i},i\in[n]$, and a repair matrix $\mathbf{M}\in\mathcal{M}_{i}$, the smaller $|\mathcal{R}(\mathbf{M})|$ is, the smaller the repair bandwidth of $C_{i}$ is. We denote the repair matrix in $\mathcal{M}_{i}$ that minimizes $\mathcal{R}(\mathbf{M})$ under the order $\preceq$ as $\mathbf{F}_{i}$, i.e.,
[…]
@@
$\bar{\beta}(\mathcal{C})=\frac{1}{n}\sum_{i\in[n]}\{BW(\mathbf{F}_{i})\}.$
@@
Note that, for every node $C_{i}$, $i\in[n]$, we always have $i\in\mathcal{R}(\mathbf{F}_{i})$.
@@
For any two distinct nodes $C_{i}$ and $C_{j},i,j\in[n]$, if $\mathcal{R}(\mathbf{F}_{i})\prec\mathcal{R}(\mathbf{F}_{j})$, then $j\notin\mathcal{R}(\mathbf{F}_{i})$.
@@
We prove this by contradiction. Suppose that $j\in\mathcal{R}(\mathbf{F}_{i})$. When $\mathcal{R}(\mathbf{F}_{i})=\mathcal{R}(\mathbf{F}_{i})$ (or $\mathcal{R}(\mathbf{F}_{i})=\mathcal{N}(\mathbf{F}_{i}^{\prime})$), we have $\mathbf{F}_{i}\in\mathcal{M}_{j}$ (or $\mathbf{F}^{\prime}_{i}\in\mathcal{M}^{\prime}_{j}$). In either case, this implies that $\mathcal{R}(\mathbf{F}_{j})\preceq\mathcal{R}(\mathbf{F}_{i})$ since $\mathcal{R}(\mathbf{F}_{j})$ (or $\mathcal{N}(\mathbf{F}_{j}^{\prime})$) is minimized under $\prec$. This contradicts the assumption that $\mathcal{R}(\mathbf{F}_{i})\prec\mathcal{R}(\mathbf{F}_{j})$. Therefore, we conclude that $j\notin\mathcal{R}(\mathbf{F}_{i})$.
Without loss of generality, we assume the code $\mathcal{C}$ satisfies $\mathcal{R}(\mathbf{F}_{1})\preceq\mathcal{R}(\mathbf{F}_{2})\preceq\dots\preceq\mathcal{R}(\mathbf{F}_{n})$. Therefore, there must exist an integer $i\in[k]$ such that $\mathcal{R}(\mathbf{F}_{i})\prec\mathcal{R}(\mathbf{F}_{k+1})$. Recall that $i\in\mathcal{R}(\mathbf{F}_{i})$ for any $i\in[n]$. If $\mathcal{R}(\mathbf{F}_{1})=\mathcal{R}(\mathbf{F}_{2})=\dots=\mathcal{R}(\mathbf{F}_{k+1})$, then $k+1]\subseteq\mathcal{R}(\mathbf{F}_{1})$, and $|\mathcal{R}(\mathbf{F}_{i})|\geq|\mathcal{R}(\mathbf{F}_{1})|\geq k+1$ for any $i\in[n]$. It follows that the repair bandwidth and repair I/O for any failed node are larger than $2k$, the trivial upper bound. Then we define $t$ as follows:
$t=\max\{i\in[k]:\mathcal{R}(\mathbf{F}_{i})\prec\mathcal{R}(\mathbf{F}_{k+1})\}.$ (12)
@@
The integer $t$ satisfies $n-\left|\mathcal{R}(\mathbf{F}_{k+1})\right|-1\leq t$ and $\bigcup_{j\in[t]}\mathcal{R}(\mathbf{F}_{j})=[t]$.
@@
By the definition of $t$, we have
$\mathcal{R}(\mathbf{F}_{t})\prec\mathcal{R}(\mathbf{F}_{t+1})=\dots=\mathcal{R}(\mathbf{F}_{k+1}).$
In particular, for any $i\in[t]$ and $j\geq t+1$, we have $\mathcal{R}(\mathbf{F}_{i})\prec\mathcal{R}(\mathbf{F}_{j})$. Combining this with Lemma 7, we get
$j\notin\bigcup_{i\in[t]}\mathcal{R}(\mathbf{F}_{i}),~\forall~j\geq t+1.$ (13)
Recall that $i\in\mathcal{R}(\mathbf{F}_{i})$ for any $i\in[n]$. It follows that
$\left\{t+1,\dots,k+1\right\}\subseteq\mathcal{R}(\mathbf{F}_{k+1}),$ (14)
and
$[t]\subseteq\bigcup_{i\in[t]}\mathcal{R}(\mathbf{F}_{i}).$ (15)
By (14), we have
$n-1-t=\left|\left\{t+1,\dots,k+1\right\}\right|\leq\left|\mathcal{R}(\mathbf{F}_{k+1})\right|.$
By (13) and (15), we can get
$\bigcup_{j\in[t]}\mathcal{R}(\mathbf{F}_{j})=[t].$
Now, we can fix the parity-check matrix of the code $\mathcal{C}$ in a systematic form as follows:
$H=\begin{bmatrix}\mathbf{A}_{1}&\dots&\mathbf{A}_{k}&\mathbf{I}_{2}&\\ \mathbf{B}_{1}&\dots&\mathbf{B}_{k}&&\mathbf{I}_{2}\end{bmatrix}$,
where all the block $\mathbf{A}_{i},\mathbf{B}_{i}$ are $2\times 2$ matrices. Further, by the MDS property, all $\mathbf{A}_{i}$ and $\mathbf{B}_{i},i\in[k]$ are invertible matrices.
By (13), we have $k+1\notin\bigcup_{i\in[t]}\mathcal{R}(\mathbf{F}_{i})$ and $k+2\notin\bigcup_{i\in[t]}\mathcal{R}(\mathbf{F}_{i})$. For any $i\in[t]$ and $j\in\{k+1,k+2\}$, $\textnormal{rank}\left(\mathbf{M}_{i}\mathbf{H}_{j}\right)\leq 1$ is guaranteed by the definition of $\mathcal{R}(\mathbf{F}_{i})$. Combining with $d(\mathbf{F}_{i})=k+1$, it implies that when we repair the node $C_{i},i\in[t]$, each of the two nodes $C_{k+1}$ and $C_{k+2}$ only needs to transmit one symbol.
From the above discussion, during the repairing of node $C_{i},i\in[t]$, we assume that the nodes $C_{n-1}$ and $C_{n}$ transmit $\mathbf{u}_{i}C_{n-1}$ and $\mathbf{v}_{i}C_{n}$, respectively, for some nonzero vectors $\mathbf{u}_{i},\mathbf{v}_{i}\in\mathbb{F}_{q}^{2}$.
@@
For each $i\in[t]$, the repair matrix $\mathbf{F}_{i}$ can be written as
$\mathbf{F}_{i}=\Lambda\begin{bmatrix}\mathbf{u}_{i}&\\ &\mathbf{v}_{i}\end{bmatrix}$,
where $\Lambda$ is a $2\times 2$ invertible matrix.
@@
Recall that node $C_{n-1}$ transmits $\mathbf{u}_{i}C_{n-1}$ and node $C_{n}$ transmits $\mathbf{v}_{i}C_{n}$ during the repair of node $C_{i}$. It follows that
$\mathbf{F}_{i}\mathbf{H}_{n-1}=\begin{bmatrix}a\mathbf{u}_{i}\\ b\mathbf{u}_{i}\end{bmatrix}\text{~and~}\mathbf{F}_{i}\mathbf{H}_{n}=\begin{bmatrix}c\mathbf{v}_{i}\\ d\mathbf{v}_{i}\end{bmatrix}$
for some $a,b,c,d\in\mathbb{F}_{q}$. Since $\mathbf{H}_{n-1}=\begin{bmatrix}\mathbf{I}_{2}\\ \mathbf{O}\end{bmatrix}$ and $\mathbf{H}_{n}=\begin{bmatrix}\mathbf{O}\\ \mathbf{I}_{2}\end{bmatrix}$, we have
$\mathbf{F}_{i}$ $=\mathbf{F}_{i}\begin{bmatrix}\mathbf{I}_{2}&\mathbf{O}\\ \mathbf{O}&\mathbf{I}_{2}\end{bmatrix}=\mathbf{F}_{i}\begin{bmatrix}\mathbf{H}_{n-1}&\mathbf{H}_{n}\end{bmatrix}$
$=\begin{bmatrix}\mathbf{F}_{i}\mathbf{H}_{n-1}&\mathbf{F}_{i}\mathbf{H}_{n}\end{bmatrix}=\begin{bmatrix}a\mathbf{u}_{i}&c\mathbf{v}_{i}\\ b\mathbf{u}_{i}&d\mathbf{v}_{i}\end{bmatrix}=\Lambda\begin{bmatrix}\mathbf{u}_{i}&\\ &\mathbf{v}_{i}\end{bmatrix}$,
where $\Lambda=\begin{bmatrix}a&c\\ b&d\end{bmatrix}$. Because the matrix $\mathbf{F}_{i}$ is of full rank, $\Lambda$ is invertible. ∎
Since the two repair matrices $\mathbf{F}_{i}$ and $\Lambda^{-1}\mathbf{F}_{i}$ behave the same during the repair process, we always assume that
$\mathbf{F}_{i}=\begin{bmatrix}\mathbf{u}_{i}&\\ &\mathbf{v}_{i}\end{bmatrix}\text{~for~all~}i\in[t].$
Now we denote the complement set of $\mathcal{R}(\mathbf{F}_{i})$ as $\overline{\mathcal{R}(\mathbf{F}_{i})}=[n]\setminus\mathcal{R}(\mathbf{F}_{i})$ for $i\in[t]$, and it is evident that $i\notin\overline{\mathcal{R}(\mathbf{F}_{i})}$. By definition, we have
$\textnormal{rank}\begin{bmatrix}\mathbf{u}_{i}\mathbf{A}_{j}\\ \mathbf{v}_{i}\mathbf{B}_{j}\end{bmatrix}=1$ (16)
for any $j\in\overline{\mathcal{R}(\mathbf{F}_{i})}\cap[k]$.

### v2

@@ 3.1 The Lower Bound on Avg-Min Repair Bandwidth
In this subsection, since we focus on the repair bandwidth, and blockwise invertible column transformation does not change the repair bandwidth (according to (3)), it is convenient to further simplify the parity-check matrix to
$H=\begin{bmatrix}I_{2}&\cdots&I_{2}&I_{2}&O\\ W_{1}&\cdots&W_{k}&O&I_{2}\end{bmatrix}$,
where each $W_{i}=B_{i}A_{i}^{-1}$ is an invertible matrix over $\mathbb{F}_{q}^{2\times 2}$.
Recall the definitions of set $\mathcal{R}(M)$ in (5) and set $\mathcal{M}_{i}$ in (6). According to $d(M)=k+1$, we have
$\operatorname{BW}(M)=\sum_{j\in[n]}\textnormal{rank}\left(MH_{j}\right)-2=k+|\mathcal{R}(M)|,$
which means that for a node $C_{i},i\in[n]$, and a repair matrix $M\in\mathcal{M}_{i}$, the smaller $|\mathcal{R}(M)|$ is, the smaller the repair bandwidth of $C_{i}$ is.
For each node $C_{i}$, let $\mathcal{R}_{i}$ be the minimum set in $\{\mathcal{R}(M):M\in\mathcal{M}_{i}\}$ under the total order $\prec$. Let $M_{i}\in\mathcal{M}_{i}$ be a repair matrix such that $\mathcal{R}(M_{i})=\mathcal{R}_{i}$. Then we have
  • $M_{i}$ can repair $C_{i}$ ($i\in\mathcal{R}_{i}$), and
  • $M_{i}$ minimizes the repair bandwidth of $C_{i}$ ($\beta_{i}=k+|\mathcal{R}_{i}|$).
@@
$\bar{\beta}(\mathcal{C})=\frac{1}{n}\sum_{i\in[n]}BW(M_{i})=k+\frac{1}{n}\sum_{i\in[n]}|\mathcal{R}_{i}|.$ (11)
@@
For every node $C_{i}$, $i\in[n]$, we have $i\in\mathcal{R}_{i}$.
@@
For any two distinct nodes $C_{i}$ and $C_{j},i,j\in[n]$, if $\mathcal{R}_{i}\prec\mathcal{R}_{j}$, then $j\notin\mathcal{R}_{i}$.
@@
Suppose, for the sake of contradiction, that $j\in\mathcal{R}_{i}$. This implies that the matrix $M_{i}$ is a valid repair matrix for node $C_{j}$ ( i.e., $M_{i}\in\mathcal{M}_{j}$). Since $\mathcal{R}_{j}$ is minimized under $\prec$, we must have $\mathcal{R}_{j}\preceq\mathcal{R}_{i}$. This contradicts the assumption that $\mathcal{R}_{i}\prec\mathcal{R}_{j}$. Therefore, we conclude that $j\notin\mathcal{R}_{i}$. ∎
For the rest of this section, we will work under the assumption that $\mathcal{R}_{1}\preceq\mathcal{R}_{2}\preceq\dots\preceq\mathcal{R}_{n}$.
@@
There exists an integer $i\in[k]$ such that $\mathcal{R}_{i}\prec\mathcal{R}_{k+1}$. Let $t:=\max\{i\in[k]:\mathcal{R}_{i}\prec\mathcal{R}_{k+1}\}.$ Then, $|\mathcal{R}_{k+1}|\geq k+1-t$.
@@
First, we verify the existence of such an integer $i$. Suppose, for the sake of contradiction, that no such $i$ exists. This would imply $\mathcal{R}_{1}=\dots=\mathcal{R}_{k+1}$. Such a condition would result in a repair bandwidth strictly greater than $2k$, contradicting Lemma 6. Thus, the set $\{i\in[k]:\mathcal{R}_{i}\prec\mathcal{R}_{k+1}\}$ is non-empty, and $t$ is well-defined.
By the definition of $t$, we have $\mathcal{R}_{t+1}=\dots=\mathcal{R}_{k+1}$. Remark 1 implies that $\{t+1,\dots,k+1\}\subseteq\mathcal{R}_{k+1}.$ Consequently, $|\mathcal{R}_{k+1}|\geq|\{t+1,\dots,k+1\}|=k+1-t$. ∎
The following lemma demonstrates the structure of the repair matrices for the first $t$ nodes.
@@
For each $i\in[t]$, the repair matrix $M_{i}$ can be decomposed as
$M_{i}=\Lambda\begin{bmatrix}\bm{u}_{i}&\\ &\bm{v}_{i}\end{bmatrix}$,
where $\Lambda\in\mathbb{F}_{q}^{2\times 2}$ is an invertible matrix and $\bm{u}_{i},\bm{v}_{i}\in\mathbb{F}_{q}^{1\times 2}$ are nonzero vectors.
@@
Since $\mathcal{R}_{i}\preceq\mathcal{R}_{t}\prec\mathcal{R}_{k+1}\preceq\mathcal{R}_{k+2}$ for each $i\in[t]$, Lemma 7 implies that $k+1,k+2\notin\mathcal{R}_{i}$.
By the definition of $\mathcal{R}_{i}$, we have $\textnormal{rank}\left(M_{i}H_{j}\right)\leq 1$ for $j\in\{k+1,k+2\}$. On the other hand, since the repair degree is $d(M_{i})=k+1$, the contribution from each helper must be nonzero, implying $\textnormal{rank}\left(M_{i}H_{j}\right)=1$ for $j\in\{k+1,k+2\}$.
Consequently, the sub-matrices $M_{i}H_{k+1}$ and $M_{i}H_{k+2}$ can be expressed as rank-1 matrices constructed from row vectors $\bm{u}_{i},\bm{v}_{i}\in\mathbb{F}_{q}^{1\times 2}$. Specifically, there exist scalars $a,b,c,d\in\mathbb{F}_{q}$ such that $M_{i}H_{k+1}=\begin{bmatrix}a\bm{u}_{i}\\ b\bm{u}_{i}\end{bmatrix}$ and $M_{i}H_{k+2}=\begin{bmatrix}c\bm{v}_{i}\\ d\bm{v}_{i}\end{bmatrix}$, where $\bm{u}_{i}$ and $\bm{v}_{i}$ are nonzero.
Note that the matrix $H_{k+1}~H_{k+2}]$ forms the identity matrix $I_{4}$. We can reconstruct $M_{i}$ as follows:
$M_{i}=M_{i}[H_{k+1}~H_{k+2}]=\Lambda\begin{bmatrix}\bm{u}_{i}&\\ &\bm{v}_{i}\end{bmatrix},\text{~where~}\Lambda=\begin{bmatrix}a&c\\ b&d\end{bmatrix}$.
Finally, since $M_{i}$ has rank $2$ and the block-diagonal matrix containing $\bm{u}_{i},\bm{v}_{i}$ has rank $2$, the matrix $\Lambda$ must also be invertible. ∎